# Remove that Square Root: A New Efficient Scale-Invariant Version of AdaGrad

### v1

@@ 3.1 Assumptions
This assumption is standard in the literature of adaptive methods [Li and Orabona, 2019, Ward et al., 2020, Liu et al., 2022]. Moreover, we assume that at any iteration $t$ of KATE, we can access $g_{t}$ — a noisy and unbiased estimate of $\nabla f(w_{t})$. We also make the following assumption on the noise of the gradient estimate $g_{t}$.
@@
$\mathbb{E}_{t}\left[\left\|g_{t}-\nabla f(w_{t})\right\|^{2}\right]\leq\sigma^{2}.$ (BV)
@@ 3.2 Main Results
$\min_{t\leq T}\left\|\nabla f(w_{t})\right\|^{2}$$\leq$ $\frac{\left(\frac{2(f(w_{0})-f_{*})}{\sqrt{\eta_{0}}\beta}+\sum_{k=1}^{d}b_{0}[k]\right)^{2}}{T+1},$
@@
Theorem 3.3 establishes an $\mathcal{O}\left(\nicefrac{{1}}{{T}}\right)$ convergence rate for KATE, which is optimal for finding a first-order stationary point of a non-convex problem [Carmon et al., 2020]. However, this result is not parameter-free. To prove the convergence, we assume that $\nu_{0}[k]\leq\frac{1}{L},\;\forall k\in[d]$ in Theorem 3.3, which is equivalent to $\beta\sqrt{1+\eta_{0}\left(\nabla_{k}f(w_{0})\right)^{2}}\leq\nicefrac{{\left(\nabla_{k}f(w_{0})\right)^{2}}}{{L}},\;\forall k\in[d]$. Note that the later condition holds for sufficiently small (dependent on $L$) values of $\beta,\eta_{0}>0$.
However, it is possible to derive a parameter-free version of Theorem 3.3. Indeed, Lemma 2.2 implies that the step sizes are decreasing. Therefore, we can break down the analysis of KATE into two phases: Phase I when $\nu_{0}[k]>\nicefrac{{1}}{{L}}$ and Phase II when $\nu_{0}[k]\leq\nicefrac{{1}}{{L}}$, when the current analysis works, and then follow the proof techniques of Ward et al. [2020] and Xie et al. [2020]. We leave this extension as a possible future direction of our work.
@@
$\mathcal{C}_{f}$$\coloneqq$ $f(w_{0})-f_{*}+2\beta\sigma\sum_{k=1}^{d}\sqrt{\eta[k]}\log\left(\frac{e(\sigma^{2}+\gamma^{2})T}{g_{0}^{2}[k]}\right)$
$+\sum_{k=1}^{d}\left(\frac{\beta^{2}\eta[k]L}{2}+\frac{\beta^{2}L}{2g_{0}^{2}[k]}\right)\log\left(\frac{e(\sigma^{2}+\gamma^{2})T}{g_{0}^{2}[k]}\right).$
@@
Theorem 3.4 shows an $\mathcal{O}\left(\nicefrac{{\log^{\nicefrac{{1}}{{2}}}T}}{{T^{\nicefrac{{1}}{{4}}}}}\right)$ convergence rate for KATE with respect to the metric $\min_{t\leq T}\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|\right]$ for the stochastic setting. Note that, in the stochastic setting, KATE achieves a slower rate than Theorem 3.3 due to noise accumulation. Up to the logarithmic factor, this rate is optimal [Arjevani et al., 2023]. Similar rates for the same metric follow from the results [^7] Défossez et al. [2020] derive ${\cal O}(\nicefrac{{\log T}}{{\sqrt{T}}})$ convergence rates for AdaGrad and Adam in terms of $\min_{t\leq T}\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|^{2}\right]$ which is not smaller than $\min_{t\leq T}\left(\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|\right]\right)^{2}$. of [Défossez et al., 2020] for AdaGrad and Adam.
Finally, Li and Orabona [2019] considers a variant of AdaGrad closely related to KATE:
$w_{t+1}=w_{t}-\frac{\beta g_{t}}{\left(\text{ diag}\left(\Delta I+\sum_{\tau=1}^{t-1}g_{\tau}g_{\tau}^{\top}\right)\right)^{\frac{1}{2}+\varepsilon}},$ (15)
for some $\varepsilon\in[0,\nicefrac{{1}}{{2}})$ and $\Delta>0$. It differs from AdaGrad in two key aspects: the denominator of the stepsize does not contain the last stochastic gradient, and also, instead of the square root of the sum of squared gradients, this sum is taken in the power of $\nicefrac{{1}}{{2}}+\varepsilon$. However, the results from Li and Orabona [2019] do not imply convergence for the case of $\varepsilon$, which is expected since, in this case, the stepsize converges to zero too quickly in general. To compensate for such a rapid decrease, in KATE, we introduce an increasing sequence $m_{t}$ in the numerator of the stepsize.
@@
(a) Plot of $f(w_{t})$
(b) Plot of Accuracy
(c) Plot of Gradient Norm
Figure 1: Comparison of KATE on scaled and un-scaled data. Figures 0(a), and 0(b) plot the functional value $f(w_{t})$ and accuracy on scaled and unscaled data, respectively. Figure 0(c) plots $\left\|\nabla f(w_{t})\right\|^{2}$ and $\left\|\nabla f(w_{t})\right\|_{V^{-2}}^{2}$ for unscaled and scaled data respectively.
@@ 4 Numerical Experiments
In this section, we implement KATE in several machine learning tasks to evaluate its performance. The code to reproduce our results can be found at https://github.com/nazya/KATE.
@@ 4.1 Logistic Regression
to elaborate on the scale-invariance and robustness of KATE for various initializations.
Synthetic data.
@@
$y_{i}=\begin{cases}1&x_{i}^{\top}Vw^{*}\geq 0,\\ -1&x_{i}^{\top}Vw^{*}<0\end{cases}\qquad\forall i\in[n].$
4.1.1 Scale Invariance
In this experiment, we implement ${\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}\sf{\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}KATE}}$ on problems (4) (for unscaled data) and (5) (for scaled data) with $\varphi_{i}(z)=\log\left(1+\exp{\left(-y_{i}z\right)}\right)$. We run ${\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}\sf{\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}KATE}}$ for $10,000$ iterations with mini-batch size $10$, $\eta=\nicefrac{{1}}{{\left(\nabla f(w_{0})\right)^{2}}}$ and plot functional value $f(w_{t})$ and accuracy in Figures 0(a) and 0(b). We use the proportion of correctly classified samples to compute accuracy, i.e. $\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\left\{y_{i}x_{i}^{\top}w_{t}\geq 0\right\}}$. In plots 0(a) and 0(b), the functional value and accuracy of ${\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}\sf{\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}KATE}}$ coincide, which aligns with our theoretical findings (Proposition 2.1). Figure 0(c) plots $\left\|\nabla f(w_{t})\right\|^{2}$ and $\left\|\nabla f(w_{t})\right\|_{V^{-2}}^{2}$ for unscaled and scaled data respectively. Here, (9) explains the identical values taken by the corresponding gradient norms of KATE iterates for the scaled and unscaled data.
4.1.2 Robustness of KATE
We compare KATE’s performance with four other algorithms: AdaGrad, AdaGradNorm, SGD-decay and SGD-constant, similar to the section 5.1 of Ward et al. [2020]. For each algorithm, we initialize with $w_{0}=0\in\mathbb{R}^{20}$ and independently draw a sample of mini-batch size $10$ to update the weight vector $w_{t}$. We compare the algorithms $\bullet$ AdaGrad with stepsize $\frac{\beta}{\sqrt{\Delta+\sum_{\tau=0}^{t}g^{2}_{\tau}}}$, $\bullet$ AdaGradNorm with step size $\frac{\beta}{\sqrt{\Delta+\sum_{\tau=0}^{t}\left\|g_{\tau}\right\|^{2}}}$, $\bullet$ SGD-decay with stepsize $\nicefrac{{\beta}}{{\Delta\sqrt{t+1}}}$, and $\bullet$ SGD-constant with step size $\nicefrac{{\beta}}{{\Delta}}$. Similarly, for KATE we use stepsize $\frac{\beta m_{t}}{b_{t}^{2}}$ where $m_{t}^{2}=\eta b_{t}^{2}+\sum_{\tau=0}^{t}\nicefrac{{g_{\tau}^{2}}}{{b_{\tau}^{2}}}$ and $b_{t}^{2}=\Delta+\sum_{\tau=0}^{t}g_{\tau}^{2}$. Here, we choose $\beta=f(w_{0})-f(w^{*})$ and vary $\Delta$ in $\{10^{-8},10^{-6},10^{-4},10^{-2},1,10^{2},10^{4},10^{6},10^{8}\}$.
@@ 4.1.2 Robustness of KATE
(a) Dataset: heart
(b) Dataset: australian
(c) Dataset: splice
@@
Figure 3: Comparison of KATE with AdaGrad, AdaGradNorm, SGD-decay and SGD-constant on datasets heart, australian, and splice from LIBSVM. Figures 2(a), 2(b) and 2(c) plot the functional value $f(w_{t})$, while 2(d), 2(e) and 2(f) plot the accuracy on $y$-axis for $5,000$ iterations.
@@ 4.1.3 Peformance of KATE on Real Data
In this section, we examine KATE’s performance on real data. We test KATE on three datasets: heart, australian, and splice from the LIBSVM library. The response variables $y_{i}$ of each of these datasets contain two classes, and we use them for binary classification tasks using a logistic regression model (16). We take $\eta=\nicefrac{{1}}{{\left(\nabla f(w_{0})\right)^{2}}}$ for KATE and tune $\beta$ in all the experiments. For tuning $\beta$, we do a grid search on the list $\{10^{-10},10^{-8},10^{-6},10^{-4},10^{-2},1\}$. Similarly, we tune stepsizes for other algorithms. We take $5$ trials for each of these algorithms and plot the mean of their trajectories.
We plot the functional value $f(w_{t})$ (i.e. loss function) in Figures 2(a), 2(b) and 2(c), whereas Figures 2(d), 2(e) and 2(f) plot the corresponding accuracy of the weight vector $w_{t}$ on the $y$-axis for $5,000$ iterations. We observe that KATE performs superior to all other algorithms, even on real datasets.
Figure 4: CIFAR10: $\eta=0$
Figure 5: CIFAR10: $\eta=0.001$
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 7: Emotion: $\eta=0$
Figure 8: Emotion: $\eta=0.001$
@@ 4.2 Training of Neural Networks
In this section, we compare the performance of KATE, AdaGrad and Adam on two tasks, i.e. training ResNet18 [He et al., 2016] on the CIFAR10 dataset [Krizhevsky and Hinton, 2009] and BERT [Devlin et al., 2018] fine-tuning on the emotions dataset [Saravia et al., 2018] from the Hugging Face Hub. We choose standard parameters for Adam ($\beta_{1}=0.9$ and $\beta_{2}=0.999$) that are default values in PyTorch and select the learning rate of $10^{-5}$ for all considered methods. We run KATE with different values of $\eta\in\{0,10^{-1},10^{-2}\}$. For the image classification task, we normalize the images (similar to Horváth and Richtárik [2020]) and use a mini-batch size of 500. For the BERT fine-tuning, we use a mini-batch size 160 for all methods.
Figures 6-9 report the evolution of top-1 accuracy and cross-entropy loss (on the $y$-axis) calculated on the test data. For the image classification task, we observe that KATE with different choices of $\eta$ outperforms Adam and AdaGrad. Finally, we also observe that KATE performs comparably to Adam on the BERT fine-tuning task and is better than AdaGrad. These preliminary results highlight the potential of KATE to be applied for training neural networks for different tasks.
@@
We organize the Supplementary Material as follows: Section A presents some technical lemmas required for our analysis. In section B, we provide the proofs for the main results of our work.

### v4

@@ 3.1 Assumptions
This assumption is standard in the literature of adaptive methods (Li and Orabona, , 2019; Ward et al., , 2020; Liu et al., , 2022; Nguyen et al., , 2018, 2021; Nguyen et al., 2017a, ; Beznosikov and Takáč, , 2021). Moreover, we assume that at any iteration $t$ of KATE, we can access $g_{t}$ — a noisy and unbiased estimate of $\nabla f(w_{t})$. We also make the following assumption on the noise of the gradient estimate $g_{t}$.
@@
$\mathbb{E}_{t}\left[\|g_{t}-\nabla f(w_{t})\|^{2}\right]\leq\sigma^{2}.$ (BV)
@@ 3.2 Main Results
$\min_{t\leq T}\left\|\nabla f(w_{t})\right\|^{2}$$\leq$ $\frac{\left(\tfrac{2(f(w_{0})-f_{*})}{\sqrt{\eta_{0}}\beta}+\sum_{k=1}^{d}b_{0}[k]\right)^{2}}{T+1},$
@@
Theorem 3.3 establishes an $\mathcal{O}\left(\nicefrac{{1}}{{T}}\right)$ convergence rate for KATE, which is optimal for finding a first-order stationary point of a non-convex problem (Carmon et al., , 2020). However, this result is not parameter-free. To prove the convergence, we assume that $\nu_{0}[k]\leq\frac{1}{L},\;\forall k\in[d]$ in Theorem 3.3, which is equivalent to $\beta\sqrt{1+\eta_{0}\left(\nabla_{k}f(w_{0})\right)^{2}}\leq\nicefrac{{\left(\nabla_{k}f(w_{0})\right)^{2}}}{{L}},\;\forall k\in[d]$. Note that the later condition holds for sufficiently small (dependent on $L$) values of $\beta,\eta_{0}>0$.
However, it is possible to derive a parameter-free version of Theorem 3.3. Indeed, Lemma 2.2 implies that the step sizes are decreasing. Therefore, we can break down the analysis of KATE into two phases: Phase I when $\nu_{0}[k]>\nicefrac{{1}}{{L}}$ and Phase II when $\nu_{0}[k]\leq\nicefrac{{1}}{{L}}$, when the current analysis works, and then follow the proof techniques of Ward et al., (2020) and Xie et al., (2020). We leave this extension as a possible future direction of our work.
@@
$\mathcal{C}_{f}$$\coloneqq$ $f(w_{0})-f_{*}+2\beta\sigma\textstyle{\sum}_{k=1}^{d}\sqrt{\eta[k]}\log\left(\tfrac{e(\sigma^{2}+\gamma^{2})T}{g_{0}^{2}[k]}\right)$
$+\textstyle{\sum}_{k=1}^{d}\left(\tfrac{\beta^{2}\eta[k]L}{2}+\tfrac{\beta^{2}L}{2g_{0}^{2}[k]}\right)\log\left(\tfrac{e(\sigma^{2}+\gamma^{2})T}{g_{0}^{2}[k]}\right).$
@@
Theorem 3.4 shows an $\mathcal{O}(\nicefrac{{\log^{\nicefrac{{1}}{{2}}}T}}{{T^{\nicefrac{{1}}{{4}}}}})$ convergence rate for KATE with respect to the metric $\min_{t\leq T}\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|\right]$ for the stochastic setting. Note that, in the stochastic setting, KATE achieves a slower rate than Theorem 3.3 due to noise accumulation. Up to the logarithmic factor, this rate is optimal (Arjevani et al., , 2023). Similar rates for the same metric follow from the results [^5] Défossez et al., (2020) derive ${\cal O}(\nicefrac{{\log T}}{{\sqrt{T}}})$ convergence rates for AdaGrad and Adam in terms of $\min_{t\leq T}\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|^{2}\right]$ which is not smaller than $\min_{t\leq T}\left(\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|\right]\right)^{2}$. of (Défossez et al., , 2020) for AdaGrad and Adam.
Finally, Li and Orabona, (2019) considers a variant of AdaGrad closely related to KATE:
$w_{t+1}=w_{t}-\frac{\beta g_{t}}{\left(\operatorname{diag}\left(\Delta I+\textstyle{\sum}_{\tau=1}^{t-1}g_{\tau}g_{\tau}^{\top}\right)\right)^{\tfrac{1}{2}+\varepsilon}},$ (16)
for some $\varepsilon\in[0,\nicefrac{{1}}{{2}})$ and $\Delta>0$. It differs from AdaGrad in two key aspects: the denominator of the stepsize does not contain the last stochastic gradient, and also, instead of the square root of the sum of squared gradients, this sum is taken in the power of $\nicefrac{{1}}{{2}}+\varepsilon$. However, the results from Li and Orabona, (2019) do not imply convergence for the case of $\varepsilon=\nicefrac{{1}}{{2}}$, which is expected since, in this case, the stepsize converges to zero too quickly in general. To compensate for such a rapid decrease, in KATE, we introduce an increasing sequence $m_{t}$ in the numerator of the stepsize.
@@ 4 Numerical Experiments
In this section, we implement KATE in several machine learning tasks to evaluate its performance. To ensure transparency and facilitate reproducibility, we provide an access to the source code for all of our experiments at https://github.com/nazya/KATE.
@@ 4.1 Logistic Regression
to elaborate on the scale-invariance and robustness of KATE for various initializations. For the experiments of this Section 4.1, we used Mac mini (M1, 2020), RAM 8 GB and storage 256 GB. Each of these plots took about 20 minutes to run.
4.1.1 Robustness of KATE
@@
$y_{i}=\left\{\begin{matrix}1,&\quad x_{i}^{\top}Vw^{*}\geq 0,\\ -1,&\quad x_{i}^{\top}Vw^{*}<0,\end{matrix}\right.\qquad\forall i\in[n].$
We compare KATE’s performance with four other algorithms: AdaGrad, AdaGradNorm, SGD-decay and SGD-constant, similar to the section 5.1 of Ward et al., (2020). For each algorithm, we initialize with $w_{0}=0\in\mathbb{R}^{20}$ and independently draw a sample of mini-batch size $10$ to update the weight vector $w_{t}$. We compare the algorithms $\bullet$ AdaGrad with stepsize $\tfrac{\beta}{\sqrt{\Delta+\sum_{\tau=0}^{t}g^{2}_{\tau}}}$, $\bullet$ AdaGradNorm with step size $\frac{\beta}{\sqrt{\Delta+\sum_{\tau=0}^{t}\left\|g_{\tau}\right\|^{2}}}$, $\bullet$ SGD-decay with stepsize $\nicefrac{{\beta}}{{\Delta\sqrt{t+1}}}$, and $\bullet$ SGD-constant with step size $\nicefrac{{\beta}}{{\Delta}}$. Similarly, for KATE we use stepsize $\frac{\beta m_{t}}{b_{t}^{2}}$ where $m_{t}^{2}=\eta b_{t}^{2}+\sum_{\tau=0}^{t}\nicefrac{{g_{\tau}^{2}}}{{b_{\tau}^{2}}}$ and $b_{t}^{2}=\Delta+\sum_{\tau=0}^{t}g_{\tau}^{2}$. Here, we choose $\beta=f(w_{0})-f(w^{*})$ and vary $\Delta$ in $\{10^{-8},10^{-6},10^{-4},10^{-2},1,10^{2},10^{4},10^{6},10^{8}\}$.
@@
(g) Dataset: heart
(h) Dataset: australian
(i) Dataset: splice
Figure 2: Comparison of KATE with AdaGrad, AdaGradNorm, SGD-decay and SGD-constant on datasets heart, australian, and splice from LIBSVM. Figures 1(d), 1(e) and 1(f) plot the functional value $f(w_{t})$, while 1(g), 1(h) and 1(i) plot the accuracy on $y$-axis for $5,000$ iterations.
@@ 4.1.2 Peformance of KATE on Real Data
In this section, we examine KATE’s performance on real data. We test KATE on three datasets: heart, australian, and splice from the LIBSVM library (Chang and Lin, , 2011). The response variables $y_{i}$ of each of these datasets contain two classes, and we use them for binary classification tasks using a logistic regression model (17). We take $\eta=\nicefrac{{1}}{{\left(\nabla f(w_{0})\right)^{2}}}$ for KATE and tune $\beta$ in all the experiments. For tuning $\beta$, we do a grid search on the list $\{10^{-10},10^{-8},10^{-6},10^{-4},10^{-2},1\}$. Similarly, we tune stepsizes for other algorithms. We take $5$ trials for each of these algorithms and plot the mean of their trajectories.
We plot the functional value $f(w_{t})$ (i.e. loss function) in Figures 1(d), 1(e) and 1(f), whereas Figures 1(g), 1(h) and 1(i) plot the corresponding accuracy of the weight vector $w_{t}$ on the $y$-axis for $5,000$ iterations. We observe that KATE performs superior to all other algorithms, even on real datasets.
@@ 4.2 Training of Neural Networks
In this section, we compare the performance of KATE, AdaGrad and Adam on two tasks, i.e. training ResNet18 (He et al., , 2016) on the CIFAR10 dataset (Krizhevsky and Hinton, , 2009) and BERT (Devlin et al., , 2018) fine-tuning on the emotions dataset (Saravia et al., , 2018) from the Hugging Face Hub. We use internal cluster with the following hardware: AMD EPYC 7552 48-Core Processor, 512GiB RAM, NVIDIA A100 40GB GPU, 200gb user storage space.
General comparison.
We choose standard parameters for Adam ($\beta_{1}=0.9$ and $\beta_{2}=0.999$) that are default values in PyTorch and select the learning rate of $10^{-5}$ for all considered methods. We run KATE with different values of $\eta\in\{0,10^{-1},10^{-2}\}$. For the image classification task, we normalize the images (similar to Horváth and Richtárik, (2020)) and use a mini-batch size of 500. For the BERT fine-tuning, we use a mini-batch size 160 for all methods.
Figures 5-8 report the evolution of top-1 accuracy and cross-entropy loss (on the $y$-axis) calculated on the test data. For the image classification task, we observe that KATE with different choices of $\eta$ outperforms Adam and AdaGrad. Finally, we also observe that KATE performs comparably to Adam on the BERT fine-tuning task and is better than AdaGrad. These preliminary results highlight the potential of KATE to be applied for training neural networks for different tasks. For BERT each run takes about 35 minutes, and 25 minutes for ResNet.
Figure 9: Cifar10: $\eta=0.001$
Figure 10: Emotion: $\eta=0.001$
Hyper-parameters tuning.
Next, we compare baselines presented in Saravia et al., (2018) for emotions classification and Zhang et al., (2019) for image classification. These papers provide efficient setups for learning rates and learning rate schedulers that are reasonable to compare with. Saravia et al., (2018) performs a search of efficient learning rate and uses a linear learning rate scheduler with warmup for Adam optimizer. A different learning rate (1e-5), $\Delta$=1e-5 and the same scheduler applied for KATE lead to the same performance, see Figure 10. We would like to point out that it is challenging to find a reference for hyper-parameters for a certain setup. Thus, to fairly compare with Saravia et al., (2018) we use distilroberta-base model. Zhang et al., (2019) did a grid search for an efficient learning rate and used a multi-step scheduler for Adam optimizer, decaying the learning rate by a factor of 5 at the 60th, 120th, and 160th epochs. Zhang et al., (2019) refers to DeVries and Taylor, (2017) for the code implementing special techniques, namely data augmentation and cutout to achieve higher accuracy. A different learning rate (1e-3), the same scheduler and $\Delta$=1e-3 applied for KATE demonstrates comparable performance, see Figure 10. For BERT each run takes about 20 minutes, while 100 minutes for ResNet.
Acknowledgments and Disclosure of Funding
We thank Francesco Orabona and Dmitry Kamzolov for the pointers to the related works that we missed while preparing the first version of this paper. We also thank anonymous reviewers for their useful feedback and suggestions.